OpenStudy (anonymous):

Add and simplify : A+2/3 + a+2/4 Answers given : A. a+2/6 B. 2a+4/12 C. 7a+14/12

OpenStudy (callisto):

Is this your question?\[\frac{a+2}{3} +\frac{a+2}{4}\]

OpenStudy (callisto):

I assume it is a yes then :| \[\frac{a+2}{3} +\frac{a+2}{4} = \frac{4(a+2)}{4(3)} +\frac{3(a+2)}{3(4)}=\frac{4a+8}{12} +\frac{3a+6}{12}\]\[=\frac{(4a+8)+(3a+6)}{12} = \frac{(4a+3a+8+6)}{12} = \frac{7a+14}{12}\]